OpenStudy (anonymous):

Differentiate the function. F(t) = ln((4t + 1)^4)/((4t − 1)^5) Will fan and medal!!!

OpenStudy (anonymous):

\[f(t)=\ln((4t+1)^4)\div ((4t-1)^5)\]

OpenStudy (loser66):

first off, apply ln (a/b) = lna - ln b then take derivative term by term.

OpenStudy (anonymous):

like this, ln(4t+1)^4-ln(4-1)^5 @Loser66

OpenStudy (loser66):

yes, then ln a^2 = 2lna , apply that rule to take down the exponent,

OpenStudy (anonymous):

3ln16-4ln20 ????@Loser66

OpenStudy (loser66):

I don't know what you mean, take second step \[ln(4t+1)^4 = 4ln(4t+1)\] then take derivative of this term do the same with the second term.

OpenStudy (loser66):

oh yes, since derivative of this term is 16/(4t+1) and the second one is 20/ (4t+1) so, the answer should be -4/(4t+1)